# RIVAL: Reinforcement Learning with Iterative and Adversarial Optimization for Machine Translation

### v1

@@ 3.2 Why Vanilla RLHF Fails in Colloquial Subtitle Translation?
Upon examination of the results, the RM achieves a high accuracy rate of $99.5\%$, attributed to the homogeneity of the data distribution. Subsequently, we apply this RM to GRPO training. Despite continuous improvement in the reward curve and no significant abnormalities in translation length, the evaluation results fall short compared to models trained via SFT using identical source texts. We even observe reward hacking phenomena where high reward scores correspond with low translation quality. For example, the model translates non-existing original texts to cheat the RM into assigning high scores.
@@ 5.2 Main Results
RIVAL-Iter0 2.69 3.4 3.32 3.48 3.26 65.22
RIVAL-Iter1 3.28 3.86 3.84 3.89 3.68 66.27
RIVAL-Iter2 3.06 3.82 3.57 3.79 3.53 66.49
@@
RIVAL-Iter0 2.48 3.15 3.10 3.26 2.96 52.41
RIVAL-Iter1 2.46 3.23 3.12 3.29 2.99 53.42
RIVAL-Iter2 2.33 3.16 3.06 3.21 2.93 51.76
Table 1: Performance comparison on in-domain translation direction (Subtitle ZH-EN) and OOD translation task (Medical ZH-DE) using GPT-4o scores and COMETKiwi. Only the qualitative preference reward is used.
@@
RIVAL-Iter0 34.76 69.12 29.52 70.16
RIVAL-Iter1 38.62 69.72 32.90 72.37
RIVAL-Iter2 39.39 72.60 33.42 73.61
@@
RIVAL-Iter0 26.44 71.50 35.35 73.49
RIVAL-Iter1 25.70 69.93 35.32 72.66
RIVAL-Iter2 25.25 68.74 35.25 72.23
Table 2: Performance comparison on in-domain WMT(EN-ZH) and OOD WMT(EN-DE) using BLEU and COMETKiwi. Both preference rewards are used.
@@ Appendix A Subtitle Case
{CJK} UTF8gbsn Source: 你将一个视频中的多个语音识别文本逐条翻译成英文。输入为一个json格式,key为序号,value为待翻译的语音文本,一共有10个文本。示例如下: 输入: {"1": "xxx", "2": "xxx"} 输出: {"1": "xxx", "2": "xxx"} 现在请根据上述要求完成如下片段的翻译,输出一共10个翻译后的结果,需要和输入一一对应。直接输出翻译后的英文,不要进行任何解释。输入: {"1": "牛老师惊恐的说:你大的要来了?我点点头:要来了,我瞒不住了。", "2": "其实我是一名鬼杀队队员。牛老师大怒:你是不是有病?我都这样了,你还不放过我?", "3": "而且还要 COS 鬼灭之刃。我叹了口气:You KNOW NOTHING ,牛牛桑,鬼灭之刃其实不是虚构的,而是一部纪录片。", "4": "世界上是真的有鬼,也是真的有鬼杀队员。", "5": "如果今天不是我及时出现,大爷们就要白发人送黑发人了。", "6": "刘老师说:那问题是我刚染了黄毛呀,不是黑发人呀。","7": "我说:我说的是你的脸,印堂发黑去非洲怎么着也得当一个王妃。", "8": "之所以你认为这个世界上没有鬼存在,都是我们这些鬼杀队队员的功劳,把你这些普通人类保护的太好。", "9": "当然,也正因为我们跟鬼残酷的战斗了几百年,以至于人才凋落,很多鬼都管不过来,才让你被鬼白嫖。", "10": "牛老师说:你快拉倒吧,你这人全身上下只有嘴硬,就你还杀鬼,你只会鬼吹灯。"} 输出: Strong Translation: {"1": "Teacher Niu said in horror: Your big one is coming? I nodded: It’s coming, I can’t hide it anymore.", "2": "Actually, I am a member of the Demon Slayer Corps. Teacher Niu was furious: Are you sick? I’m already like this, and you still won’t let me go?", "3": "And I have to COS Demon Slayer. I sighed: You KNOW NOTHING, Niu Niu-san, Demon Slayer is not fictional, it’s actually a documentary.", "4": "There really are demons in the world, and there really are demon slayers.", "5": "If I hadn’t appeared in time today, the old men would have had to send off the young.", "6": "Teacher Liu said: But the problem is I just dyed my hair yellow, I’m not a young person with black hair.", "7": "I said: I’m talking about your face, with a darkened forehead, you could at least be a princess in Africa.", "8": "The reason you think there are no demons in this world is all thanks to us demon slayers, protecting you ordinary humans too well.", "9": "Of course, it’s also because we’ve been fighting brutally with demons for hundreds of years, leading to a decline in talent, and many demons are left unchecked, which is why you got taken advantage of by demons.", "10": "Teacher Niu said: Stop it, you’re all talk and no action, you killing demons? You only know how to blow hot air."}" {CJK} UTF8gbsn Source: "请你将一个视频中的多个语音识别文本逐条翻译成英文。输入为一个json格式,key为序号,value为待翻译的语音文本,一共有10个文本。示例如下: 输入: {"1": "xxx", "2": "xxx"} 输出: {"1": "xxx", "2": "xxx"} 现在请根据上述要求完成如下片段的翻译,输出一共10个翻译后的结果,需要和输入一一对应。直接输出翻译后的英文,不要进行任何解释。 输入: {"1": "啊哈。但是你听到后面跟的是个降调的感觉。听啊,", "2": "啊,再来一下。", "3": "当然了,倒水最主要声音最响就是这个叭叭这种声音。", "4": "贯穿始终有一种大珠小珠落玉盘的感觉,它呢,其实就是这种入水的声音。就是这,", "5": "是吧?", "6": "我应该找 B .Boxer 酋长来录这一期的。", "7": "总之呢,倒一杯水有三种声音诶,嘚嘚吧嘚吧。", "8": "但是在科学上,你不能用得儿劲儿蹦来描述谁的声音。听着不太聪明的样子。", "9": "声音的本质是物体的振动。那这三种声音都是啥在振动呢?那个声调的怎?", "10": "N 是杯子里面水面上方空气柱的振动。也就是我们小学二年级学过的海姆霍兹共振。"} 输出: " Strong Translation: "{"1": "Aha. But you hear that the following part has a falling tone. Listen,", "2": "Ah, let’s do it again.", "3": "Of course, the loudest sound when pouring water is this kind of ’ba ba’ sound.", "4": "Throughout, there is a feeling of big and small beads falling on a jade plate. It is actually the sound of water entering. It’s this,", "5": "Right?", "6": "I should have invited B.Boxer Chief to record this episode.", "7": "In short, there are three kinds of sounds when pouring a glass of water, de de ba de ba.", "8": "But scientifically, you can’t describe someone’s sound as ’de er jin er beng’. It doesn’t sound very smart.,̈ "9": T̈he essence of sound is the vibration of objects. So what are these three sounds vibrating? What about the pitch?", "10": "N is the vibration of the air column above the water surface inside the cup. This is the Helmholtz resonance we learned in second grade."}"
@@ Appendix B GPT-4o Evaluation Prompt
{CJK} UTF8gbsn GPT-4o Evaluation Prompt 如下: 请你以公正的评判者的身份,结合参考译文,对翻译文本的效果进行评估。请保持一致的评估标准。你需要按照下面的维度对翻译文本进行评估,根据每个维度的要求描述,给每个维度打出一个14̃的分数。 维度1:准确性:词句+专有名词翻译是否准确 1分:翻译严重偏离原义,3处以上词语存在事实性错误或语义矛盾。 2分:翻译部分偏离原义,2-3处词语存在事实性错误或语义矛盾。 3分:1处词语存在事实性错误或语义矛盾,或是无词语翻译错误,仅有语义错误。 4分:译文完全准确,无事实性错误,且不影响用户感知。 补充说明: 1. 如果翻译存在漏翻,如果未导致句子主要含义发生变化,则只扣完整性,不扣准确性分数; 2. 原文的错别字、断句不需要结合上下文进行纠正,直接按照原文文本翻译即可,翻译正确即不扣准确性分数。 3. 增添部分逻辑合理,不做准确性扣分。 维度2:完整性:是否漏翻关键词句 1分:句子主要信息和逻辑关系严重缺失,主从句不完整,数词、动词及动词受者大量缺失,达3处以上遗漏问题,严重影响原义理解。 2分:句子主要信息部分缺失主句和从句(非状语、非补语)部分未保留主谓语,关键成分(数词、动词及其受者)未完整保留,达2-3处遗漏问题,不影响整体理解。 3分:句子主要信息部分缺失,主句和从句(非状语、非补语)部分未保留主谓语,关键成分(数词、动词及其受者)仅小部分遗漏,仅1处遗漏问题,不影响整体理解。 4分:内容完整性:确保句子的主要信息和逻辑关系100%完整。主从句完整性:主句和从句(非状语、非补语)必须保留主谓语。仅允许进行浓缩或代词替换。其他成分处理:宾语、补语、插入语和修饰语可以进行压缩性意译或删除。关键成分保留:数词、动词及其受者不可省略,必须完整保留。 补充说明: 1)请严格对比原文,看主要信息是否均翻译完成,可以接受对语义重复的修饰语做删除,不扣完整性分数; 2)原文的错别字、断句错误导致的问题,不扣完整性分数。 维度3:连贯性:上下文衔接是否连贯 1分:语法错误频发,上下文转折生硬,逻辑断裂,严重影响用户理解。 2分:多处语序混乱或句式杂糅,上下文衔接略生硬,关联词使用不准确,需反复理解 3分:整体连贯,仅存在部分的生硬衔接,存在过长修饰语或出现少量语句重复。 4分:语句衔接自然,表达顺畅,符合口语表达习惯。
{CJK} UTF8gbsn 补充说明: 1)连词、独立语、过渡性语句等的省略,如果对句子内容产生影响,则只扣完整性分数,连贯性不扣分; 2)因为原文错别字、断句导致的问题,不对此维度扣分。 维度4:风格一致性:上下文叙述风格是否一致 1分:翻译内容与视频语境不符,3处以上语法结构、用词与视频风格不一致,使用过于突兀。 2分:翻译内容在书面语和口语之间切换,2-3处用词和句法结构不符合语境。 3分:上下文风格大体一致,自然流畅,仅1处用词和句法结构不符合语境。 4分:上下文内容呈现出统一的口语或书面语的特征,整体内容风格明显。 补充说明: 首先判断视频文本的场景为书面还是口语场景,后判断语法结构、用词是否符合场景要求。 我们会提供原文、参考译文、待评估的翻译文本,你需要依照前面的每个维度标准,分别给出翻译文本在各个维度的打分并解释理由。请注意,评分应尽可能严格。如果样例在某个维度上同时符合低分和高分的评价标准,应优先选择低分,而非较高的分数。只有当样例完全符合高分标准时,才能给予高分。你的回答以 JSON 格式输出结果,严格按照如下格式返回: {"维度1": {"解释": "解释", "打分": 1}} 【以下是原文】 {origin_text} 【以下是参考译文】 {reference_text} 【以下是待评估的翻译文本】 {translate_text}
@@ Appendix E MAE vs MSE case
{CJK}
UTF8gbsn Source: 请将以下内容翻译为中文: 235. At the first regular session of 1999 of the Administrative Committee on Coordination, in April, the members concluded that, to meet the challenges of globalization, the United Nations system needed to cooperate more effectively with the private sector and civil society, as well as with Governments.

### v2

@@ 3.2 Why Vanilla RLHF Fails in Colloquial Subtitle Translation?
[…]
@@ 5.2 Main Results
RIVAL-Iter0-Qual 2.69 3.4 3.32 3.48 3.26 65.22
RIVAL-Iter1-Qual 3.28 3.86 3.84 3.89 3.68 66.27
RIVAL-Iter2-Qual 3.06 3.82 3.57 3.79 3.53 66.49
@@
RIVAL-Iter0-Qual 2.48 3.15 3.10 3.26 2.96 52.41
RIVAL-Iter1-Qual 2.46 3.23 3.12 3.29 2.99 53.42
RIVAL-Iter2-Qual 2.33 3.16 3.06 3.21 2.93 51.76
Table 1: Performance comparison on in-domain translation direction (Subtitle ZH-EN) and OOD translation task (Medical ZH-DE) using GPT-4o scores and COMETKiwi. Only the qualitative preference reward is used. All results are averaged by 3 times.
@@
RIVAL-Iter0-Qual 33.34 69.64 28.32 69.91
RIVAL-Iter1-Qual 31.74 70.83 27.65 71.19
RIVAL-Iter2-Qual 30.14 71.91 26.73 73.28
RIVAL-Iter0-Qual+Quant 34.76 69.12 29.52 70.16
RIVAL-Iter1-Qual+Quant 38.62 69.72 32.90 72.37
RIVAL-Iter2-Qual+Quant 39.39 72.60 33.42 73.61
@@
RIVAL-Iter0-Qual 25.15 71.48 35.17 73.69
RIVAL-Iter1-Qual 24.71 70.93 34.97 73.06
RIVAL-Iter2-Qual 23.96 69.81 34.82 72.76
RIVAL-Iter0-Qual+Quant 26.44 71.50 35.35 73.49
RIVAL-Iter1-Qual+Quant 25.70 69.93 35.32 72.66
RIVAL-Iter2-Qual+Quant 25.25 68.74 35.25 72.23
Table 2: Performance comparison on in-domain WMT(EN-ZH) and OOD WMT(EN-DE) using BLEU and COMETKiwi. Both preference rewards are used. All results are averaged by 3 times.
@@
Different reward signals serve different purposes. As shown in Table 2, our RIVAL method achieves comparable performance under both qualitative and quantitative rewards. However, when the quantitative reward is incorporated, the corresponding BLEU score improves significantly. In fact, omitting the quantitative reward leads to a drop in BLEU but an increase in COMET, suggesting that the model is able to explore effectively at the semantic level without being constrained by surface-level lexical similarity.
@@
In addition, we found that using only qualitative rewards can still maintain better performance than the SFT method OOD scenarios. Furthermore, on both BLEU and COMETkiwi metrics, it shows a similar pattern to the in-domain scenario: training with qualitative rewards leads to a significant decrease in character-level similarity but not at the semantic level.
@@ Appendix A Subtitle Case
Strong Translation: {"1": "Teacher Niu said in horror: Your big one is coming? I nodded: It’s coming, I can’t hide it anymore.", "2": "Actually, I am a member of the Demon Slayer Corps. Teacher Niu was furious: Are you sick? I’m already like this, and you still won’t let me go?", "3": "And I have to COS Demon Slayer. I sighed: You KNOW NOTHING, Niu Niu-san, Demon Slayer is not fictional, it’s actually a documentary.", "4": "There really are demons in the world, and there really are demon slayers.", "5": "If I hadn’t appeared in time today, the old men would have had to send off the young.", "6": "Teacher Liu said: But the problem is I just dyed my hair yellow, I’m not a young person with black hair.", "7": "I said: I’m talking about your face, with a darkened forehead, you could at least be a princess in Africa.", "8": "The reason you think there are no demons in this world is all thanks to us demon slayers, protecting you ordinary humans too well.", "9": "Of course, it’s also because we’ve been fighting brutally with demons for hundreds of years, leading to a decline in talent, and many demons are left unchecked, which is why you got taken advantage of by demons.", "10": "Teacher Niu said: Stop it, you’re all talk and no action, you killing demons? You only know how to blow hot air."}" Source: "请你将一个视频中的多个语音识别文本逐条翻译成英文。输入为一个json格式,key为序号,value为待翻译的语音文本,一共有10个文本。示例如下:
输入: {"1": "xxx", "2": "xxx"}
输出: {"1": "xxx", "2": "xxx"}
现在请根据上述要求完成如下片段的翻译,输出一共10个翻译后的结果,需要和输入一一对应。直接输出翻译后的英文,不要进行任何解释。
输入: {"1": "啊哈。但是你听到后面跟的是个降调的感觉。听啊,", "2": "啊,再来一下。", "3": "当然了,倒水最主要声音最响就是这个叭叭这种声音。", "4": "贯穿始终有一种大珠小珠落玉盘的感觉,它呢,其实就是这种入水的声音。就是这,", "5": "是吧?", "6": "我应该找 B .Boxer 酋长来录这一期的。", "7": "总之呢,倒一杯水有三种声音诶,嘚嘚吧嘚吧。", "8": "但是在科学上,你不能用得儿劲儿蹦来描述谁的声音。听着不太聪明的样子。", "9": "声音的本质是物体的振动。那这三种声音都是啥在振动呢?那个声调的怎?", "10": "N 是杯子里面水面上方空气柱的振动。也就是我们小学二年级学过的海姆霍兹共振。"}
输出: "
Strong Translation: "{"1": "Aha. But you hear that the following part has a falling tone. Listen,", "2": "Ah, let’s do it again.", "3": "Of course, the loudest sound when pouring water is this kind of ’ba ba’ sound.", "4": "Throughout, there is a feeling of big and small beads falling on a jade plate. It is actually the sound of water entering. It’s this,", "5": "Right?", "6": "I should have invited B.Boxer Chief to record this episode.", "7": "In short, there are three kinds of sounds when pouring a glass of water, de de ba de ba.", "8": "But scientifically, you can’t describe someone’s sound as ’de er jin er beng’. It doesn’t sound very smart.,̈ "9": T̈he essence of sound is the vibration of objects. So what are these three sounds vibrating? What about the pitch?", "10": "N is the vibration of the air column above the water surface inside the cup. This is the Helmholtz resonance we learned in second grade."}"
@@ Appendix B GPT-4o Evaluation Prompt
GPT-4o Evaluation Prompt 如下: 请你以公正的评判者的身份,结合参考译文,对翻译文本的效果进行评估。请保持一致的评估标准。你需要按照下面的维度对翻译文本进行评估,根据每个维度的要求描述,给每个维度打出一个14̃的分数。 维度1:准确性:词句+专有名词翻译是否准确 1分:翻译严重偏离原义,3处以上词语存在事实性错误或语义矛盾。 2分:翻译部分偏离原义,2-3处词语存在事实性错误或语义矛盾。 3分:1处词语存在事实性错误或语义矛盾,或是无词语翻译错误,仅有语义错误。 4分:译文完全准确,无事实性错误,且不影响用户感知。 补充说明: 1. 如果翻译存在漏翻,如果未导致句子主要含义发生变化,则只扣完整性,不扣准确性分数; 2. 原文的错别字、断句不需要结合上下文进行纠正,直接按照原文文本翻译即可,翻译正确即不扣准确性分数。 3. 增添部分逻辑合理,不做准确性扣分。 维度2:完整性:是否漏翻关键词句 1分:句子主要信息和逻辑关系严重缺失,主从句不完整,数词、动词及动词受者大量缺失,达3处以上遗漏问题,严重影响原义理解。 2分:句子主要信息部分缺失主句和从句(非状语、非补语)部分未保留主谓语,关键成分(数词、动词及其受者)未完整保留,达2-3处遗漏问题,不影响整体理解。 3分:句子主要信息部分缺失,主句和从句(非状语、非补语)部分未保留主谓语,关键成分(数词、动词及其受者)仅小部分遗漏,仅1处遗漏问题,不影响整体理解。 4分:内容完整性:确保句子的主要信息和逻辑关系100%完整。主从句完整性:主句和从句(非状语、非补语)必须保留主谓语。仅允许进行浓缩或代词替换。其他成分处理:宾语、补语、插入语和修饰语可以进行压缩性意译或删除。关键成分保留:数词、动词及其受者不可省略,必须完整保留。 补充说明: 1)请严格对比原文,看主要信息是否均翻译完成,可以接受对语义重复的修饰语做删除,不扣完整性分数; 2)原文的错别字、断句错误导致的问题,不扣完整性分数。 维度3:连贯性:上下文衔接是否连贯 1分:语法错误频发,上下文转折生硬,逻辑断裂,严重影响用户理解。 2分:多处语序混乱或句式杂糅,上下文衔接略生硬,关联词使用不准确,需反复理解 3分:整体连贯,仅存在部分的生硬衔接,存在过长修饰语或出现少量语句重复。 4分:语句衔接自然,表达顺畅,符合口语表达习惯。
补充说明: 1)连词、独立语、过渡性语句等的省略,如果对句子内容产生影响,则只扣完整性分数,连贯性不扣分; 2)因为原文错别字、断句导致的问题,不对此维度扣分。 维度4:风格一致性:上下文叙述风格是否一致 1分:翻译内容与视频语境不符,3处以上语法结构、用词与视频风格不一致,使用过于突兀。 2分:翻译内容在书面语和口语之间切换,2-3处用词和句法结构不符合语境。 3分:上下文风格大体一致,自然流畅,仅1处用词和句法结构不符合语境。 4分:上下文内容呈现出统一的口语或书面语的特征,整体内容风格明显。 补充说明: 首先判断视频文本的场景为书面还是口语场景,后判断语法结构、用词是否符合场景要求。 我们会提供原文、参考译文、待评估的翻译文本,你需要依照前面的每个维度标准,分别给出翻译文本在各个维度的打分并解释理由。请注意,评分应尽可能严格。如果样例在某个维度上同时符合低分和高分的评价标准,应优先选择低分,而非较高的分数。只有当样例完全符合高分标准时,才能给予高分。你的回答以 JSON 格式输出结果,严格按照如下格式返回: {"维度1": {"解释": "解释", "打分": 1}} 【以下是原文】 {origin_text} 【以下是参考译文】 {reference_text} 【以下是待评估的翻译文本】 {translate_text}
@@ Appendix E MAE vs MSE case
Source: 请将以下内容翻译为中文: 235. At the first regular session of 1999 of the Administrative Committee on Coordination, in April, the members concluded that, to meet the challenges of globalization, the United Nations system needed to cooperate more effectively with the private sector and civil society, as well as with Governments.